OpenStudy (anonymous):

18x+12y=8 30x+24y=14 find x and y. it says use an inverse matrix function but is there another way to do this problem?

OpenStudy (anonymous):

there are a couple of different ways to do it

OpenStudy (anonymous):

you could use either substitution or elimination method to solve this

jimthompson5910 (jim_thompson5910):

did you still need help, or did you manage to figure it out?

OpenStudy (anonymous):

jim, you cant let me have nothing can ya lol

jimthompson5910 (jim_thompson5910):

lol sry, go ahead

OpenStudy (anonymous):

lol its all good

OpenStudy (anonymous):

thank you sooo much

OpenStudy (anonymous):

@NickR when i have problems like that cant i just use the elimination method and then plug in whichever variable i find to get the other?because i dont like using the matrix method

OpenStudy (anonymous):

yes you can

OpenStudy (anonymous):

i knew it. our teacher makes everything so comlplicated

OpenStudy (anonymous):

lol ya i dont even remember the matrix method. and the elimination method is what abhi showed you

OpenStudy (anonymous):

yes sir! are u tired?

OpenStudy (anonymous):

little bit why?

OpenStudy (anonymous):

i have another question. .2x-.6y=3.2 -x+1.4y=-11.2 i used eliminiation and got that x=21 y=7 and idk why thats wrong!

OpenStudy (anonymous):

x-3y=16 -x+1.4y=-11.2 -1.6y=4.8 y=-3

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

sorry i took so long to reply